100cm³ of oxygen that fixed through an office in 60 seconds while it take 120seconds for the same office. h=1 o=16?


1
Expert's answer
2021-05-24T07:17:46-0400

Rate of effusion of O2 = effused volume / time .

100 / 60 = V / 120 .


100 × 120 / 60 = V .


V = 200 Cm3 = 200 ml.